OpenStudy (anonymous):

8(n)=7(n-6) solving each proportions

OpenStudy (saifoo.khan):

8n = 7n - 42 8n - 7n = -42 n = -42

OpenStudy (anonymous):

okay where did you get this from

OpenStudy (saifoo.khan):

I showed you the working.. and there is NO typo.

OpenStudy (anonymous):

because i got fifty with 42+8 = 50 then 50/7=7.142857 142857 142857 142857